ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: Singh, S. N., Univ. of Transkei, Umtata 5117, South Africa Solutions of boundary value problems by double Laplace Transformations In solutions of boundary value problems, previous researchers have applied Laplace transformations with respect to one variable only. In the present paper, we have applied Laplace transformations with respect to both the variables in a partial differential equation involving functions of two variables and have obtained the solutions with greater ease.